OpenStudy (anonymous):

-16x^-1 y^-3 ------------- 80x^-4 y^-3 Simplify. rewrite the expression with positive exponents.

OpenStudy (cwrw238):

-16x^-1 y^-3 = - 16 ------------ x y^3 80x^-4 y^-3 = 80 ----------- x^4 y^3 invert the last fraction and multiply = 16 x^4 y^3 ------- * ------- xy^3 80 now simplify and thats it

OpenStudy (cwrw238):

can you continue with this?

OpenStudy (cwrw238):

oh - i missed out the '-' before the 16 on last line

OpenStudy (anonymous):

I will be honest...I have no clue how to do it from there.

OpenStudy (cwrw238):

well 16divides into 80 5 times , the y^3's cancel out and x^4 / x = x^3